OpenStudy (anonymous):

PLEASE HELP! MEDAL WILL BE REWARDED!!! Find The Derivative g(x)=^4(sqroot)x

OpenStudy (ddcamp):

Is that: \[\sqrt[4]{x}\ or\ 4\sqrt{x}\]

OpenStudy (anonymous):

First one

OpenStudy (ddcamp):

Radicals can be rewritten as fractional exponents: \[\sqrt[n]{x} \rightarrow x^{\frac{ 1 }{ n }}\] Once the expression is in exponent form, you can use the power rule.

OpenStudy (anonymous):

so would I get 1/4x^3/4

OpenStudy (ddcamp):

Not quite. What's \[\frac{ 1 }{ 4 } - 1?\]

OpenStudy (anonymous):

Its -3/4

OpenStudy (ddcamp):

So \[\frac{1}{4}x^{\frac{-3}{4}}\] Is your final answer

OpenStudy (anonymous):

Oh wow thank you so much!!! If I post another problem would you help? Ill give you another medal!

OpenStudy (ddcamp):

OK